OpenStudy (anonymous):

i just need someone to check my differential equation \[D^2y + 2D y + 50y = 0; \; y(0) = 7; \; y'(0) = 7\] i got \[y= y_c = e^{-x} (7\cos 7x + 2\sin 7x)\]

OpenStudy (dominusscholae):

Correct.

OpenStudy (unklerhaukus):

\[\Delta=b^2-4ac=2^2-4\times1\times50=-196\]\[m=\frac{-b\pm\sqrt{\Delta}}{2a}\]\[=\frac{-2\pm\sqrt{-196}}{2}\]\[m=-1\pm7i\] \[y_c=e^{-x}(B\cos 7x+C\sin7x)\]

OpenStudy (anonymous):

so...im right right?

OpenStudy (unklerhaukus):

\[y(0)=7\] \[y(0)=e^{-0}(B\cos (7\cdot0)+C\sin(7\cdot0))\]\[7=B\cos (7\cdot0)\]\[7=B\]

OpenStudy (anonymous):

uh-huh...

OpenStudy (anonymous):

yes u are :)

OpenStudy (unklerhaukus):

\[y=e^{-x}(7\cos 7x+C\sin7x)\]\[y'=e^{-x}(-49\sin 7x+7C\cos7x) -e^{-x}(7\cos 7x+C\sin7x)\]\[y'=-e^{-x}((C+49)\sin 7x+7(C+1)\cos7x)\] \[\qquad y'(0)=7\] \[y'(0)=-e^{-0}((C+49)\sin (7\cdot0)+7(C+1)\cos(7\cdot 0))\]\[7=-7(C+1)\]?

OpenStudy (anonymous):

negative/ o.O

OpenStudy (unklerhaukus):

i might have made a mistake